OpenStudy (anonymous):

What is the equation of the line that passes through (1, 3) and (-2, -3)?

OpenStudy (mew55):

1.find the slope rise/run (y2-y1)/(x2-x1)= -3-3/-2-1=-6/-3=3 then with this equation y-y0=m(x-x0) y-3=2(x-1) y-3=2x-2 y=2x+1 i chose the first set of points, since it will be easier.